# The role of turbulence in setting the phase of the ISM and implications for the star formation rate

### v1

@@ 4.5 Caveats of the simulations
Figure 9: Relations of the SFR with column density for different gas reservoirs. In the left panel, the contours sketch the location of the data point from Bigiel et al. (2008): the outer contour tracers roughly all points, while inner one corresponds roughly to their orange contours denoting 2 independent data point per 0.05 dex. In the middle panel, we show a power-law with index 1.4 to guide the eye. In the right panel, we show the linear relation found by Bigiel et al. (2008) between the SFR and the column density of H2 gas, and extrapolate to to lower densities. The simulation data is colour-coded as in Figure 6.
@@ 5.1 SFR versus different gas phases
As discussed in the introduction, observations have revealed correlations between the gas surface density and the SFR. In Figure 9 we correlate the SFR with different gas reservoirs in the simulations. The SFR is measured over time intervals of 0.1 Myr. Observations usually distinguish between HI and H2, the later estimated from CO emission. Our simulations do not track chemical evolution and it is non-trivial to estimate the fraction of the neutral gas which would be H2. Instead, we will simply separate the dense gas component, that is cold gas with densities exceeding $n_{\mathrm{neutral}}>100$ cm-3 as in Section 4.3.
@@
In the left panel of Figure 9, we show the total neutral gas versus the SFR. We compare with the observational results from Bigiel et al. (2008), who studied the KS-relation in a sample of 18 galaxies. Their results for HI + H2 are shown as contours, where the outer contour sketches the full extend of their data points and the inner contour indicate where most of the measurements are situated. The observations are characterised by a break at 9 $\mathrm{M}_{\odot}$pc-2. For higher column densities, a power-law relation with index 1.4 is found, while for lower column densities the relation is steeper with a lot of scatter (Kennicutt & Evans, 2012). The results from the high-$\Sigma$ runs fall in the observed range, though the SFR might be a bit high compared to the majority of the observations. Possibly this is because we do not include early feedback in the form of stellar winds and jets, which are expected to reduce the SFR by a factor 2 (Rathjen et al., 2021; Verliat et al., 2022). At low column density, we reproduce the large scatter. The simulations without driving produce SFRs which are below the observed range. This could mean that the environment modelled by this simulation is unlikely to occur in real galaxies. Indeed, large-scale dynamic effects such as rotation or infall would provide a minimal level of turbulence driving. The simulations with strongest driving (orange and green points) generally agree best with the observations.
@@
The middle panel shows the relation with cold gas, that is CNM and cold dense gas which would be molecular. A positive correlation now appears at low column densities. These findings are in line with the results from Smith et al. (2023) who studied the distribution of CNM in their full galaxy simulation. Overall, the relation between the column density of cold gas and the SFR exhibits a scatter of about one order of magnitude around a power-law with index 1.4. This is the same index as found for the relation with total gas, above a column density of 20 $\mathrm{M}_{\odot}$pc-2. This leads us to the conclusion that the scatter in the observations of the KS-relation at low column densities, as shown in the left panel, is likely due to regional differences in the CNM fraction. Following this line of thought, it implies that at higher column densities the CNM fraction on a scale of 1 kpc does not vary significantly between regions. Indeed, in the simulations we find it to be around 40 % independent of the external driving strength.
@@ 5.2 Knee-density of the KS-relation: two regimes
The kink in the KS-relation is observed around 9 $\mathrm{M}_{\odot}$pc2, in between the low- and high-$\Sigma$ runs. The findings from our simulations suggest this marks a transition between two regimes of star formation, each with a different dominant limiting process.
@@
From our simulations, it is not clear whether turbulence still plays an important role in the creation of cold clouds at high densities. We saw that strong external driving keeps a significant amount of the gas in the thermally unstable phase. Stellar feedback also injects turbulence and so a certain minimum level of turbulence driving can be associated to a SFR. Together with the additional UV heating of young stars, this may lead to a self-regulation of the CNM-fraction on kpc scales in regions with $\Sigma>$ 9 $\mathrm{M}_{\odot}$pc2.
@@ 6 Model for turbulence induced CNM formation
The simulation results indicate that turbulence can trigger the cooling of gas initially in the WNM state, and that this may be an important regulator for star formation in the low density regions. The mechanism behind this is similar as for the balance of turbulence with gravity: turbulence creates overdensities. Since cooling is efficient in high density gas, overdensities created by shocks cool fast to the CNM state.
@@ 6.5 Parameter dependence of the predicted CNM fraction
$f_{\mathrm{CNM}}=\frac{1}{N}\int_{n_{\mathrm{thr}}(G_{0})}^{\infty}\mathcal{P}(n,n_{0},\sigma)n\mathrm{d}n,$ (6)
where $N$ is a normalisation factor such that $0\leq f_{\mathrm{CNM}}\leq 1$,
$N=\int_{0}^{\infty}\mathcal{P}n\mathrm{d}n.$ (7)
The resulting CNM fraction and its dependence on $\sigma$, $n_{0}$ and $G_{0}$ is illustrated in Figure 12.
The behaviour of $f_{\mathrm{CNM}}$ with increasing $\sigma$ depends on how the average density compares with the CNM threshold density. If $n_{0}<n_{\mathrm{thr}}$, the CNM fraction is zero at low levels of turbulence and increases steadily with $\sigma$. When $n_{0}\approx n_{\mathrm{thr}}$, we still find an increasing behaviour of $f_{\mathrm{CNM}}$ with $\sigma$ but a point of diminishing returns is reached where increasing the turbulence further provides a smaller gain in $f_{\mathrm{CNM}}$. This is because the PDF broadens not only to high densities but also to low densities as large voids are being created as counterparts to the dense shocks. Once $n_{0}>n_{\mathrm{thr}}$, the medium is predicted to be fully in the CNM state at low levels of turbulence. Adding turbulence now typically broadens the density PDF to densities below the CNM threshold, slightly decreasing $f_{\mathrm{CNM}}$ in favour of the creation of LNM and WNM. This is a theoretical form of CNM destruction by turbulence which was not clearly observed in the simulations.
@@ 6.6 Comparison to simulation results
For each snapshot, we measure the average total density of neutral gas $n_{0}$ and the velocity dispersion $\sigma$ in the mid-plane region to input into our model for $f_{\mathrm{CNM}}$. The average density varies over time as the vertical balance in the disk first evolves from the initial conditions to balance the various supporting pressure forces against gravity, and later on adapts to the varying levels of star formation activity. We find values between 0.35 and 1.55 cm-3 across simulations. The velocity dispersion results from the combination of external turbulence driving on large scales and stellar feedback. We find values between 2 and 25 km s-1. We set $G_{0}=0.6$, as this is the value of the constant UV background used in the cooling routines.
@@ 6.6.1 Gas phase fractions
In Figure 13, we compare the model predicted phase fraction to the ones measured in the simulations [^5] Note that the measured CNM fractions here are higher than those shown in Figure 6 since here we are limiting ourselves to the densest part of the disk.. While our main goal is to model the amount of CNM, we also verify what the model predicts for the LNM. For the low-$\Sigma$ runs, which have fairly low $f_{\mathrm{CNM}}$, the prediction is excellent (left panel, bottom left corner). The model accurately describes the trend of increasing $f_{\mathrm{CNM}}$ with increasing external turbulence driving strength. The agreement is less satisfying for the high-$\Sigma$ simulations. While $f_{\mathrm{CNM}}$ is predicted correctly in the strong driving case, it is underestimated in the other three simulations. The discrepancy can be slightly reduced by increasing the parameter $b$ of Eq. 5 for these three runs. This is justified by their higher SFR which lead to more SN feedback, a mostly compressive turbulence driver. Gravitational compression is also stronger when the average density is higher.
The prediction for the LNM fraction is shown in the middle panel. In most cases, it is overestimated and for the high-$\Sigma$ runs we do not reproduce the expected increase with turbulence driving strength. Our model does not take into account the unstable nature of the LNM, which could explain why we overestimate it. Looking at the right panel, we see that the agreement between the model and the simulations is generally good for the sum $f_{\mathrm{CNM}}+f_{\mathrm{LNM}}$ (or equivalently $f_{\mathrm{WNM}}$). This indicates that once in the LNM state, the gas is likely to quickly cool down further to form CNM.
@@ 6.6.2 Density PDF
Figure 14: Comparison between the measured mass-weighted density PDF and the modelled PDF for a representative snapshot of each simulation. The black curve corresponds to the total neutral gas in the simulation, while the orange and blue curves show the contribution from the WNM and CNM respectively. The dashed line show the model calculated using the average density and velocity dispersion measured in the snapshot. Only the inner part of the disk, between -100 and +100 pc around the mid-plane, is considered for the measurement of the PDF and average properties. The low-$\Sigma$ runs are shown in the top panels, the high-$\Sigma$ ones in the bottom. External driving strength increases from left to right.
We can gain more insight in the results by directly comparing the measured density PDF with the modelled one. In Figure 14, we show such a comparison for a selection of example snapshots, one for each simulation. Generally, the WNM peak around 1 cm-3 is quite well-described by the model. The point of transition to CNM is also captured reasonably well. What does not match well is the shape of the CNM component. In most cases, the model does not reproduce the position of the peak of the distribution. This component of the PDF is also too broad, resulting in an overestimation of the amount of dense CNM. Even though the shape of the CNM component leaves much to be desired, the relative importance of CNM versus WNM is well-captured, resulting in a good prediction for $f_{\mathrm{CNM}}$.
@@ 7 Summary and conclusions
In this study, we investigate the effect of large-scale turbulence driving on star formation and the thermal balance of the ISM. We analyse a suite of numerical simulations of a 1 kpc3 part of a galaxy including stellar feedback in the form of SN and HII regions. The cooling model reproduces a multi-phase ISM with warm and cold neutral phase. The simulation suite explores the star formation behaviour for two initial column density values representative of regions in the Milky Way: 10.2 and 19.1 $\mathrm{M}_{\odot}$pc-2. For each subset of simulations, we tested four different strengths of the external large-scale turbulence driving, including a reference simulation without this external component.
We found that the additional turbulence is not only capable of decreasing the SFR, as found in previous studies, but can also enhance it under the right conditions. In the simulations with a lower column density, the SFR increases with increasing external driving strength and the star formation activity is easily an order of magnitude higher than in the case without driving. In the simulations with higher column density, we saw that when the external LS driving becomes significantly stronger than the turbulence injection from SN, the SFR is reduced.
This behaviour is linked to the thermal state of the gas. In the lower column density simulations, increasing the level of turbulence boosts CNM formation which in turn leads to more dense gas for star formation. The CNM fraction increases with driving strength. For the higher column density simulations, the CNM fraction is similar for all runs, while the LNM fraction increases with increasing levels of turbulence as it acts as a perturbing force and promotes phase change. Additionally, the amount of dense CNM is reduced, leading to a smaller star formation rate.
Studying the relation between the SFR and the column density of different mass reservoirs, we further characterise the two regimes. Overall we find a linear relation between the dense gas and the SFR, as in observations. For the lower column density runs we identify an additional correlation between the SFR and the CNM. Our interpretation is that at low densities star formation is limited by the formation of cold gas clouds. At higher densities, an abundance of cold clouds exist and the limiting process is the formation of dense clumps inside those clouds. This change in regime may be an explanation for the observed knee in the Kennicutt-Schmidt relation. This would imply that the scatter in this relation at low column densities is due to variations in the CNM fraction.
@@
Our analytical model predicts that when the average density of a region is low compared to the threshold density for CNM formation, increasing the level of turbulence will boost the CNM fraction. This is because more shocks and sharper density contrasts are created at higher Mach numbers, resulting in a widening of the density PDF which now extend to values above the CNM threshold. On the other hand, if the average density was already larger than the CNM threshold, all the gas will be in the cold phase at low levels of turbulence. Increasing the turbulent velocity dispersion broadens the PDF both to higher and lower densities, resulting in the generation of WNM and a small decrease in CNM fraction.
Using this model, we are able to predict the CNM fraction from most of the simulations. The match is especially good in the low column density regime where the WNM is the dominant phase. However, when directly comparing the model density PDF to the one measured in the simulations, we see that the shape of the CNM component is not well-described. This shows that more work is needed to accurately model the complex ISM.

### v2

@@ 5.5 Constant or variable UV background?
Figure 12: Relations of the SFR with column density for different gas reservoirs. Top: simulation set with constant UV (Table 1), bottom: set with variables UV (Table 2). The SFR is averaged over a time scale of 10 Myr after (full circles) or before (empty circles) the time of the snapshot at which the gas column densities have been measured. In the left panels, the contours sketch the location of the data point from Bigiel et al. (2008): the outer contour tracers roughly all points, while inner one corresponds roughly to their orange contours denoting 2 independent data point per 0.05 dex. In the middle panels, we show a power-law with index 1.4 to guide the eye. In the right panel, we show the linear relation found by Bigiel et al. (2008) between the SFR and the column density of H2 gas, and extrapolate to lower densities. The simulation data is colour-coded as in Figures 6 and 10.
@@ 6.1 SFR versus different gas phases
As discussed in the introduction, observations have revealed correlations between the gas surface density and the SFR. In Figure 12 we correlate the SFR with different gas reservoirs in the simulations. The SFR is measured by averaging 10 Myr after (full circles) or before (empty circles) the time of the snapshot at which the gas column densities have been measured. We verified that our conclusions do not depend on the value chosen for the time-scale of averaging by recomputing the figure for times between 2.5 Myr and 40 Myr. In fact, the correlations presented slightly improve with increasing averaging times. Observations usually distinguish between HI and H2, the later estimated from CO emission. The simulations with constant UV field do not track chemical evolution and it is non-trivial to estimate the fraction of the neutral gas which would be H2. Instead, we will simply separate the dense gas component, that is cold gas with densities exceeding $n_{\mathrm{neutral}}>100$ cm-3 as in Section 4.3. The correlation between SFR and dense gas is omitted for the simulation with variable UV background (bottom row). This is because the spatial resolution of these simulations is lower and therefore the dense gas is not as well described as in the first set of simulations.
@@
In the left panels of Figure 12, we show the total neutral gas versus the SFR. We compare with the observational results from Bigiel et al. (2008), who studied the KS-relation in a sample of 18 galaxies. Their results for HI + H2 are shown as contours, where the outer contour sketches the full extend of their data points and the inner contour indicate where most of the measurements are situated. The observations are characterised by a break at 9 $\mathrm{M}_{\odot}$pc-2. For higher column densities, a power-law relation with index 1.4 is found, while for lower column densities the relation is steeper with a lot of scatter (Kennicutt & Evans, 2012). The results from the high-$\Sigma$ runs with constant UV field (top panels) fall in the observed range, though the SFR might be a bit high compared to the majority of the observations. Possibly this is because we do not include early feedback in the form of stellar winds and jets, which are expected to reduce the SFR by a factor 2 (Rathjen et al., 2021; Verliat et al., 2022). At low column density, we reproduce the large scatter. The simulations without driving produce SFRs which are below the observed range. This could mean that the environment modelled by this simulation is unlikely to occur in real galaxies. Indeed, large-scale dynamic effects such as rotation or infall would provide a minimal level of turbulence driving. The simulations with strongest driving (orange and green points) generally agree best with the observations.
The behaviour is slightly different for the simulations with variable UV field (bottom panels). They populate additional regions of the diagram. While many points follow the observations closely, two of the simulations stand out from the others. The simulation with an initial column density of 8.4 $\mathrm{M}_{\odot}$ pc-2 without driving exhibits relatively high SFRs for low column densities of neutral gas. Such points were completely absent for the first set of simulations. They correspond to times in the second half the simulation, after a significant portion of the gas has been ionised by the radiative feedback of massive stars, leading to the low columns densities of neutral material. The other outlier is the simulation with a column density of 12.7 $\mathrm{M}_{\odot}$ pc-2 and strong driving. For this run, the SFR is lower than what is expected based on the observations during most of its evolution. This possibly indicates that such high levels of turbulence are not representative of real situations with similar column densities.
@@
The middle panels shows the relation with cold gas, that is CNM and cold dense gas which would be molecular. A positive correlation now appears at low column densities. These findings are in line with the results from Smith et al. (2023) who studied the distribution of CNM in their full galaxy simulation. Overall, the relation between the column density of cold gas and the SFR exhibits a scatter of about one order of magnitude around a power-law with index 1.4. This is the same index as found for the relation with total gas, above a column density of 20 $\mathrm{M}_{\odot}$ pc-2. This leads us to the conclusion that the scatter in the observations of the KS-relation at low column densities, as shown in the left panel, is likely due to regional differences in the CNM fraction. Following this line of thought, it implies that at higher column densities the CNM fraction on a scale of 1 kpc does not vary significantly between regions. Indeed, in the simulations we find it to be around 40 % independent of the external driving strength.
Also for the simulations with variables UV background in the bottom panel, we find a similar correlation between the SFR and the cold gas. Again, the strong driving run deviates in behaviour: there is a large variation in SFR for similar column densities. The large velocity dispersions in these runs (see Figure 11), may prevent the formation of sink-forming gas, as discussed in Section 4.3.
@@ 6.2 Knee-density of the KS-relation: two regimes
The kink in the KS-relation is observed around 9 $\mathrm{M}_{\odot}$pc-2. The findings from our simulations suggest this marks a transition between two regimes of star formation, each with a different dominant limiting process. We propose the following theory.
@@
From our simulations, it is not clear whether turbulence still plays an important role in the creation of cold clouds at high densities. We saw that strong external driving keeps a significant amount of the gas in the thermally unstable phase. Stellar feedback also injects turbulence and so a certain minimum level of turbulence driving can be associated to a SFR. Together with the additional UV heating of young stars, this may lead to a self-regulation of the CNM-fraction on kpc scales in regions with $\Sigma>$ 9 $\mathrm{M}_{\odot}$pc-2.
6.3 Current models for the break in the KS-relation
In the literature, we find two main theories for the origin of the break in the KS-relation.
The model of Ostriker et al. (2010) emphasizes the role played by local thermal and vertical mechanical equilibrium in the galactic disk. In particular, a transition in the KS relation is predicted around $\simeq 10$ $\mathrm{M}_{\odot}$ pc-2 as a consequence of the gravity being primarily dominated by the stellar potential or the self-gravity of the gas (see for instance their Figure 2).
The other idea is that the location of the knee is associated to the phase change from HI to H2, instead of from WMN to CNM as proposed in this work (Schaye, 2004; Krumholz et al., 2009b; Krumholz, 2013). This idea is supported by the data from observations which reveals that the SFR is linearly proportional to H2. It also relies on the assumption that star formation occurs in molecular gas exclusively.
An important difference between the scenario explored in this paper and these previous studies is that a relatively large set of environmental conditions is explored. As can be seen from Figure 12, this provides a broad variety of situations. Apart from the break of the mean slope around 10 M pc-2 in the KS diagram, the dispersion of the SFR at lower column densities is very large, almost 3 dex. Our work suggests that this is possibly a consequence of the 2-phase physics and the sensibility of the WNM/CNM transition to the local UV background as well as the local turbulent driving. Since our simulations are not placed in a general, self-consistent, galactic context it is however hard, at this stage, to assess to what extend the variability of the environmental conditions is sufficient to reproduce the observed variability of the KS relations.
Another important aspect that we are not addressing here is the role of the metallicity. It is now well established that metallicity has a significant role on the HI component in galaxies (Wong et al., 2013; Schruba et al., 2018) and does affect the molecular gas content. The possible role metallicity has on the SFR is a bit more controversial. Some studies (e.g. Jameson et al., 2016) found possible significant dependences of the SFR with metallicities in agreement with the predictions of the models presented by Ostriker et al. (2010) and Krumholz (2013) Other studies (Cormier et al., 2014; Bothwell et al., 2016) found that whereas metallicities has a strong influence on the HI and H2 contents, it has a weak impact on the star formation rate itself. This latter conclusion is also met by the numerical simulations of molecular clouds performed by Glover & Clark (2012) where it has been found that the star formation rate would vary by a factor of only 2 when the metallicity was changed by a factor of a hundred.
@@ 7 Model for turbulence induced CNM formation
The simulation results indicate that turbulence can trigger the cooling of gas initially in the WNM state, and that this may be an important regulator for star formation in the low density regions that are exposed to a significant UV background. The mechanism behind this is similar as for the balance of turbulence with gravity: turbulence creates overdensities. Since cooling is efficient in high density gas, overdensities created by shocks cool fast to the CNM state.
@@ 7.4 Modified density PDF of a two-phase ISM
When replacing $S$ by $S(n)$, we are not guaranteed to preserve the normalisation of the lognormal PDF. To solve this issue, we introduce two coefficients $A$ and $B$,
$\mathcal{P}(\delta)$ $=$ $\dfrac{A}{\sqrt{2\pi S(\delta)}}\exp\left(-\dfrac{(\delta-\frac{S(\delta)}{2}+B)^{2}}{2S(\delta)}\right)$ (7)
such that
$\int\mathcal{P}(\delta)\,\mathrm{d}\delta=1$ (8)
$\int e^{\delta}\mathcal{P}(\delta)\,\mathrm{d}\delta=1$ (9)
This ensures that the PDF is well normalized in probability and in mass.
@@ 7.5 Parameter dependence of the predicted CNM fraction
$f_{\mathrm{CNM}}=\int_{n_{\mathrm{thr}}(G_{0})}^{\infty}e^{\delta}\mathcal{P}(\delta,n_{0},\sigma)\mathrm{d}\delta.$ (10)
Equation (9) ensures that the result is properly normalised: $0\leq f_{\mathrm{CNM}}\leq 1$. The resulting CNM fraction and its dependence on $\sigma$, $n_{0}$ and $G_{0}$ is illustrated in Figure 15.
The behaviour of $f_{\mathrm{CNM}}$ with increasing $\sigma$ depends on how the average density compares with the CNM threshold density. If $n_{0}<n_{\mathrm{thr}}$, the CNM fraction is zero at low levels of turbulence and increases steadily with $\sigma$. When $n_{0}>n_{\mathrm{thr}}$, the medium is predicted to be fully in the CNM state at low levels of turbulence. Adding turbulence now typically broadens the density PDF to densities below the CNM threshold, slightly decreasing $f_{\mathrm{CNM}}$ in favour of the creation of LNM and WNM. This is a theoretical form of CNM destruction by turbulence which was not clearly observed in the simulations. The transition occurs around $n_{0}\approx n_{\mathrm{thr}}$, where we find a non-zero minimum in the $f_{\mathrm{CNM}}(\sigma)$ curve. At very high $\sigma$, a point of diminishing returns is reached where increasing the turbulence further provides a smaller gain in $f_{\mathrm{CNM}}$. This is because the PDF broadens not only to high densities but also to low densities as large voids are being created as counterparts to the dense shocks.
@@ 7.6 Comparison to simulation results
For each snapshot, we measure the average total density of neutral gas $n_{0}$ and the velocity dispersion $\sigma$ in the mid-plane region to input into our model for $f_{\mathrm{CNM}}$. The average density varies over time as the vertical balance in the disk first evolves from the initial conditions to balance the various supporting pressure forces against gravity, and later on adapts to the varying levels of star formation activity. We find values between 0.35 and 1.55 cm-3 across the simulations in the set with constant UV background. The velocity dispersion results from the combination of external turbulence driving on large scales and stellar feedback. We find values between 2 and 25 km s-1. We set $G_{0}=0.6$, as this is the value of the constant UV background used in the cooling routines.
@@ 7.6.1 Gas phase fractions
In Figure 16, we compare the model predicted phase fraction to the ones measured in the simulations with constant UV background [^9] Note that the measured CNM fractions here are higher than those shown in Figure 6 since here we are limiting ourselves to the densest part of the disk.. While our main goal is to model the amount of CNM, we also verify what the model predicts for the LNM. For the low-$\Sigma$ runs, which have fairly low $f_{\mathrm{CNM}}$, the prediction is excellent (left panel, bottom left corner). The model accurately describes the trend of increasing $f_{\mathrm{CNM}}$ with increasing external turbulence driving strength. The agreement is less satisfying for the high-$\Sigma$ simulations. While $f_{\mathrm{CNM}}$ is predicted correctly in the strong driving case, it is underestimated in the other three simulations. The discrepancy can be reduced by increasing the parameter $b$ of Eq. 6 for these three runs. This is justified by their higher SFR which lead to more SN feedback, a mostly compressive turbulence driver. Gravitational compression is also stronger when the average density is higher.
The prediction for the LNM fraction is shown in the middle panel. In most cases, it is slightly overestimated and for the high-$\Sigma$ runs we do not reproduce the expected increase with turbulence driving strength. Our model does not take into account the unstable nature of the LNM, which could explain why we overestimate it.
Looking at the right panel, we see that the agreement between the model and the simulations for the sum $f_{\mathrm{CNM}}+f_{\mathrm{LNM}}$ (or equivalently $f_{\mathrm{WNM}}$ is good for low column density. For the high-$\Sigma$ runs, the WNM fraction is overestimated by the model.
@@ 7.6.2 Density PDF
Figure 17: Comparison between the measured mass-weighted density PDF and the modelled PDF for a representative snapshot of each simulation. The black curve corresponds to the total neutral gas in the simulation, while the orange and blue curves show the contribution from the WNM and CNM respectively. The dashed line shows the model calculated using the average density and velocity dispersion measured in the snapshot. Only the inner part of the disk, between -100 and +100 pc around the mid-plane, is considered for the measurement of the PDF and average properties. The low-$\Sigma$ runs are shown in the top panels, the high-$\Sigma$ ones in the bottom. External driving strength increases from left to right.
We can gain more insight in the results by directly comparing the measured density PDF with the modelled one. In Figure 17, we show such a comparison for a selection of example snapshots, one for each simulation. Generally, the WNM peak around 1 cm-3 is quite well-described by the model. The point of transition to CNM is also captured reasonably well. What does not match well is the shape of the CNM component. In most cases, the model does not reproduce the position of the peak of the distribution. This component of the PDF is also too broad in some cases, resulting in an overestimation of the amount of dense CNM. Even though the shape of the CNM component leaves much to be desired, the relative importance of CNM versus WNM is well-captured, resulting in a good prediction for $f_{\mathrm{CNM}}$.
@@ 8 Summary and conclusions
In this study, we investigate the effect of large-scale turbulence driving on star formation and the thermal balance of the ISM. We analyse a suite of numerical simulations of a 1 kpc3 part of a galaxy including stellar feedback in the form of SN and HII regions. The cooling model reproduces a multi-phase ISM with warm and cold neutral phase. The simulation suite explores the star formation behaviour for several initial column density values representative of regions in the Milky Way ranging from 8.4 to 19 $\mathrm{M}_{\odot}$pc-2 . For each subset of simulations, we tested different strengths of the external large-scale turbulence driving, including a reference simulation without any external driving. Regarding the UV background, two assumptions are considered, either a constant value or a time-dependent one, proportional to the star formation rate in the computational box.
We found that the additional turbulence is not only capable of decreasing the SFR, as found in previous studies, but can also enhance it under the right conditions. In a medium with low column density exposed to a constant UV background the SFR increases with increasing external driving strength and the star formation activity is easily an order of magnitude higher than in the case without driving. When the column density was increased, we saw that when the external LS driving becomes significantly stronger than the turbulence injection from SN, the SFR is reduced. Also in the case of an SFR-dependent UV background, additional turbulence queches star formation.
The changes in SFR are linked to the thermal state of the gas. In the lower column density simulations with constant UV background, increasing the level of turbulence boosts CNM formation which in turn leads to more dense gas for star formation. The CNM fraction increases with driving strength. For the higher column density simulations, the CNM fraction is similar for all runs, while the LNM fraction increases with increasing levels of turbulence as it acts as a perturbing force and promotes phase change. Additionally, the amount of dense CNM is reduced, leading to a smaller star formation rate. In the simulations with a time-dependent UV background, the CNM fraction is reduced when additional external turbulence is present.
Studying the relation between the SFR and the column density of different mass reservoirs, we further characterise the two regimes. Overall we find a linear relation between the dense gas and the SFR, as in observations. We identify an additional correlation between the SFR and the CNM. This is particularly striking for the low column density runs with constant UV background, were no correlation is found between the SFR and the full mass reservoir of neutral gas (including WNM). Our interpretation is that at low densities or high UV backgrounds, star formation is limited by the formation of cold gas clouds. At higher densities or low UV backgrounds, an abundance of cold clouds exist and the limiting process is the formation of dense clumps inside those clouds. This change in regime may be an explanation for the observed knee in the Kennicutt-Schmidt relation. This would imply that the scatter in this relation at low column densities is due to variations in the CNM fraction.
@@
Our analytical model predicts that when the average density of a region is low compared to the threshold density for CNM formation (which is dependent on the UV background), increasing the level of turbulence will boost the CNM fraction. This is because more shocks and sharper density contrasts are created at higher Mach numbers, resulting in a widening of the density PDF which now extend to values above the CNM threshold. On the other hand, if the average density was already larger than the CNM threshold, all the gas will be in the cold phase at low levels of turbulence. Increasing the turbulent velocity dispersion broadens the PDF both to higher and lower densities, resulting in the generation of WNM and a small decrease in CNM fraction.
Using this model, we are able to predict the CNM fraction for most of the simulations with a constant UV background. The match is especially good in the low column density regime where the WNM is the dominant phase. However, when directly comparing the model density PDF to the one measured in the simulations, we see that the shape of the CNM component is not well-described. This shows that more work is needed to accurately model the complex ISM.
To conclude, in general turbulence quenches star formation. But when the heating-cooling balance is in favour of forming WNM rather than CNM, which can be the case when the region is subject to an external heating source or lacks the metals that provide the main ISM cooling, turbulence can boost star formation by providing overdensities in which the ISM can cool more easily. It remains to be investigated, using observational data, where these types of environmental conditions occur in the real ISM, and whether this can indeed be linked to the knee in the Kennicutt-Schmidt relation.